# How Well Can Knowledge Edit Methods Edit Perplexing Knowledge?

### v1

@@ 5 Experiment
Table 1: CounterFact data Pearson correlation betweeen "perplexingness" and edit efficacy (* indicates corresponding entry has p-value below 0.05).
@@
To quantify this relationship, Pearson correlation coefficients are computed and are presented in Table 1. Additionally, to assess the statistical significance of these correlations, p-values are calculated. Entries corresponding to p-values falling below the significance threshold of 0.05 are marked with * within the table.
It is observed that all the coefficients’ p-values are beneath the 0.05 threshold, thereby indicating a statistically significant correlation between "perplexingness" and edit efficacy. This means that when a model finds new knowledge very perplexing, it is difficult to incorporate this knowledge into the model. Similarly, a person might be resistant to learning something they find hard to understand.
Furthermore, the analysis reveals that certain scenarios exhibit high Pearson coefficients, such as the application of MEMIT to the GPT-2 large model. This variance could stem from the possibility that different models encode "perplexingness" in distinct manners, and that editing methods may interact with this "perplexingness" uniquely.
@@
5.2 Hierarchical relations
To enhance our understanding of the factors contributing to "perplexingness", we introduce a dataset named HierarchyData. This dataset is aimed at investigating whether hierarchical relations between words can affect "perplexingness", subsequently influencing the edit efficacy.
Significantly higher "perplexingness" of higher hierarchy level knowledge
Do hierarchical relations affect "perplexingness"? We divide the HierarchyData into two groups: hypernyms (abstract concepts) and hyponyms (specific concepts). For example, a statement like "A British Shorthair is a kind of cat" represents a specific level, while "A cat is a kind of animal" exemplifies an abstract level. To investigate the effect of hierarchical relations on "perplexingness," we analyze these two groups. The box plots are included in Appendix D. We conduct t-tests for two independent samples to determine if the mean "perplexingness" of the specific level is statistically lower than that of the abstract level. The results of the t-tests are detailed in Table 2, with all values demonstrating statistical significance. Our findings indicate that knowledge on a higher hierarchical level (more abstract) is associated with greater "perplexingness" for the models. This suggests that hierarchical relations are a factor affecting knowledge "perplexingness" for language models.
@@ 5.2 Hierarchical relations
$0.00728^{*}$ $0.00605^{*}$ $1.330e-06^{*}$
Table 2: Comparative analysis of "perplexingness" in HierarchyData: t-test results for specific vs. abstract level distributions (* indicates corresponding entry has p-value below 0.05).
Correlations between "perplexingness" and edit efficacy
Next, we aim to determine if the correlation between "perplexingness" and edit efficacy also holds for the HierarchyData dataset. We employ the same method to analyze HierarchyData as analyzing CounterFact, focusing on the Pearson correlation coefficient between "perplexingness" and edit efficacy. The right panel of Figure 3 provides one of the scatter plots (see Appendix B for other plots), showcasing the application of MEMIT on GPT2-XL. We also calculate the Pearson coefficients, with the results presented in Table 3. In this table, p-values below 0.05 are marked with *, indicating statistical significance. Our analysis reveals a consistent trend: an increase in "perplexingness" correlates with poorer efficacy of edits (higher negative log conditional probability). This pattern holds true across all scenarios, except when applying the ROME and MEMIT techniques to the GPT-J model.
Figure 3: Pre vs. post probability of new knowledge (MEMIT on GPT2-XL). a.CounterFact(left). b.HierarchyData(right).
@@
Table 3: HierarchyData Pearson correlation betweeen "perplexingness" and edit efficacy (* indicates corresponding entry has p-value below 0.05)
Relationships between hierarchical relations and edit efficacy
Additionally, we want to determine if hierarchical relations within the knowledge ultimately affect the edit efficacy. Box plots (see Appendix C) are constructed to visually compare the efficacy across the two hierarchical levels. Figure 5 shows one of the examples. Furthermore, we conduct t-tests on two independent samples to determine whether the mean of the specific level distribution is significantly lower than that of the abstract level distribution. The p-values obtained are documented in Table 4. This finding underscores a markedly lower efficacy in editing knowledge at higher hierarchical levels (more abstract knowledge). Significantly, this discrepancy indicates that hierarchical relationships profoundly affect the efficacy of specific editing techniques, like ROME and MEMIT, when applied to particular models, such as GPT2-Large and GPT2-XL. For fine-tuning and LoRA, the results do not appear to be significant, possibly because these methods can address knowledge at different hierarchical levels similarly. But, how about GPT-J?
Figure 4: Same knowledge "perplexingness" in different models (HierarchyData).
@@
From the previous experiment, we observe that GPT-J did not show any difference in edit efficacy when editing higher hierarchy and lower hierarchy knowledge. To determine if GPT-J finds the same knowledge less perplexing compared to GPT-2L and GPT-2XL, we generated a heatmap of each knowledge’s ’perplexingness’ in the HierarchyData for each model, as shown in Figure 4. Each line represents a piece of knowledge in the HierarchyData, sorted by "perplexingness" in the GPT-2L model. We observed that GPT-J appears darker in the heatmap, indicating it finds the same knowledge less perplexing.
@@
Table 4: Comparative analysis of efficacy in HierarchyData: t-test results for specific vs. abstract level distributions (* indicates corresponding entry has p-value below 0.05).
Lack of Significant Findings Across Knowledge Categories
Besides hierarchical relations, we also try to find if categories of knowledge would affect "perplexingness". We attempt to categorize the data based on the types of knowledge; however, this method does not yield any significant insights related to "perplexingness".
@@ 6 Discussion
Our experimental results reveal intriguing variations in how different models handle perplexing knowledge, particularly in the context of editing. Specifically, the application of ROME and MEMIT to GPT-J exhibits a notably low Pearson correlation between "perplexingness" and editing efficacy. Moreover, within the HierarchyData context, these correlations appear insignificant. Additionally, the influence of hierarchical relations on the editing efficacy of ROME and MEMIT when applied to GPT-J seems negligible. This suggests that GPT-J may employ a unique mechanism for storing and processing different hierarchy level knowledge compared to other models. These differences highlight the need to comprehend each model’s unique architecture and methods for handling perplexing concepts, suggesting a move towards tailored editing strategies.
@@
An intuition is that when editing towards a hypernym (“animal” $\rightarrow$ “plant”), it is assumed that the hyponym (“cat” $\rightarrow$ “plant”) is edited as well, making the edit of hypernym inherently harder. Yet, the dependent knowledge is usually not edited, for popular editing methods (Li et al., 2023).
Are there other factors that may influence the "perplexingness"?
The investigation into the responsiveness of different editing techniques to perplexing knowledge reveals that FT and LoRA are seemingly unaffected by the hierarchical structure of knowledge. Notably, there exists a pronounced correlation between "perplexingness" and the efficacy of edits. This suggests that while FT and LoRA are adept at navigating the hierarchical relationships among words, they falter when addressing the inherent "perplexingness" present within the knowledge. This observation leads to the hypothesis that additional factors, beyond hierarchical complexity, play a pivotal role in influencing "perplexingness" when employing FT and LoRA for knowledge editing.
@@
The impact of "perplexingness" on the efficacy of various editing methodologies can vary significantly. Moreover, the manner in which different models interpret, process, and encode the "perplexingness" of knowledge also differs. This suggests a complex interplay between the editing methods used and the intrinsic mechanisms of the models, underscoring the need for a nuanced understanding of both to optimize knowledge editing strategies.
@@
a. Future model editing efforts should pay attention to understanding the nature of the knowledge being edited, particularly its level of "perplexingness". To aid in this endeavor, we have introduced a hierarchy dataset designed to facilitate it. It is crucial to ensure that editing methods are versatile and effective across a diverse range of data types. b. Moreover, adopting different editing approaches tailored to the specificities of each model can significantly enhance the success of edits. And when edit hierarchy knowledge, we can try to use edit methods like fine-tune or LoRA. It may dismiss the influence of hierarchy data. c. Also, we should pay attention to the side effect of knowledge edit.
@@ 7 Conclusion
In our study, we focus on the challenges of updating large language models (LLMs) with perplexing knowledge. We meticulously define "perplexingness" and efficacy respectively. Through a comprehensive analysis using the CounterFact dataset, we identify a significant negative correlation between the "perplexingness" of the new knowledge and the efficacy of the edits across diverse scenarios. This core finding underscores the variable efficacy of editing methods in handling knowledge with different levels of initial "perplexingness".
Furthermore, we develop a specialized dataset HierarchyData, consisting of hyponym-hypernym pairs. This dataset, emphasizing hierarchical relations, serves as a tool for a more contextual evaluation of edit efficacy. We undertake a thorough review of current knowledge editing methodologies using this dataset. Our findings reveal that abstract knowledge are inherently more perplexing to LLMs than their specific counterparts. Also, our investigation into the impact of hierarchical knowledge structures on edit outcomes reveal that more abstract knowledge exhibits lower editing efficacy in some scenarios. Our methodology and dataset collectively provide a novel and rigorous approach to evaluating the efficacy of knowledge edits, offering valuable insights into the factors that contribute to their success or failure.
Our investigation into the targeted editing of knowledge within LLMs sheds light on a previously underexplored facet of model editing technology. The findings underscore the challenges associated with editing knowledge that spans various levels of "perplexingness", revealing significant discrepancies in editing efficacy. This research not only enriches our understanding of the inherent complexities in model editing but also sets a foundational basis for the development of more sophisticated editing methodologies in the future. By pushing the boundaries of our current capabilities, we move closer to achieving more refined and precise manipulations of knowledge within these advanced AI systems, marking a significant step forward in the evolution of LLMs.
@@ 8 Limitation
a. In this paper, we focus on a short hierarchy chain to facilitate the comparison between higher and lower hierarchy levels. We have not yet explored longer hierarchy chains. b. The experiment can be scaled up, including the use of larger models and larger datasets. c. Additional types of evaluation can be applied. For instance, we could ask language models specific questions to determine if the knowledge has actually been edited. However, this approach is very labor-intensive and was not implemented in this study.
Appendix A Correlation of "perplexingness" and efficacy in CounterFact
We plot the "perplexingness" (pre-edit probabilities of the new target) against the efficacy (post-edit probabilities of the new target) to visually analyze their relationship. This analysis is conducted using the first 2000 groupings from the CounterFact dataset. Figure 6 displays the scatter plot for editing methods applied to GPT2-Large. Similarly, Figure 7 presents the scatter plot for methods used on GPT2-XL, and Figure 8 illustrates the scatter plot for edits performed on GPT-J(6B).
Figure 6: Pre vs. post probability of new knowledge (CounterFact) on GPT2-Large using a. FT (upper left) b. LoRA (upper right) c. ROME (lower left) d. MEMIT (lower right).
Figure 7: Pre vs. post probability of new knowledge (CounterFact) on GPT2-XL using a. FT (upper left) b. LoRA (upper right) c. ROME (lower left) d. MEMIT (lower right).
Figure 8: Pre vs. post probability of new knowledge (CounterFact) on GPT-J(6B) using a. FT (upper left) b. LoRA (upper right) c. ROME (lower left) d. MEMIT (lower right).
Appendix B Correlation of "perplexingness" and efficacy in HierarchyData
To visually explore the relationship between "perplexingness" and editing efficacy, we plot these dimensions against each other using 198 groupings from the HierarchyData dataset. Figure 9 shows the scatter plot highlighting the effects of editing methods on the GPT2-Large model. Likewise, Figure 10 demonstrates the scatter plot for the GPT2-XL model, and Figure 11 displays the scatter plot for edits on the GPT-J(6B) model, providing a clear visual representation of how "perplexingness" correlates with the efficacy of knowledge edits across different models.
Figure 9: Pre vs. post probability of new knowledge (HierarchyData) on GPT2-Large using a. FT (upper left) b. LoRA (upper right) c. ROME (lower left) d. MEMIT (lower right).
Figure 10: Pre vs. post probability of new knowledge (HierarchyData) on GPT2-XL using a. FT (upper left) b. LoRA (upper right) c. ROME (lower left) d. MEMIT (lower right).
Figure 11: Pre vs. post probability of new knowledge (HierarchyData) on GPT-J(6B) using a. FT (upper left) b. LoRA (upper right) c. ROME (lower left) d. MEMIT (lower right).
@@ Appendix B Correlation of "perplexingness" and efficacy in HierarchyData
Figure 12: Specific vs. abstract probability distribution (HierarchyData) on GPT2-Large using a. FT (upper left) b. LoRA (upper right) c. ROME (lower left) d. MEMIT (lower right).
Figure 13: Specific vs. abstract probability distribution (HierarchyData) on GPT2-XL using a. FT (upper left) b. LoRA (upper right) c. ROME (lower left) d. MEMIT (lower right).
Figure 14: Specific vs. abstract probability distribution (HierarchyData) on GPT-J(6B) using a. FT (upper left) b. LoRA (upper right) c. ROME (lower left) d. MEMIT (lower right).
@@
We perform a comparative analysis of the "perplexingness" across both specific and abstract hierarchical levels by plotting their distributions. This analysis is based on 198 instances from the HierarchyData dataset, evenly divided between 99 specific-level and 99 abstract-level cases. Figure 15 presents the box plots, illustrating the impact of editing methods on the GPT2-Large, GPT2-XL, and GPT-J(6B) models, thereby offering insights into the variation of "perplexingness" across different levels of hierarchy and models.
Figure 15: Pre-edit specific vs. abstract probability distribution (HierarchyData) on a. GPT2-Large (upper left) b. GPT2-XL (upper right) c. GPT-J(6B) (below).

### v2

@@ 6.1 Correlations between perplexingness and edit ineffectiveness
Table 1: CounterFact data Pearson correlation between perplexingness and edit ineffectiveness ( indicates corresponding entry has $p$-value below 0.05).
@@
To quantify this relationship, Pearson correlation coefficients are computed and are presented in Table 1. Additionally, to assess the statistical significance of these correlations, $p$-values are calculated. Entries corresponding to $p$-values falling below the significance threshold of 0.05 are marked with within the table.
It is observed that all the coefficients’ $p$-values are beneath the 0.05 threshold, thereby indicating a statistically significant correlation between perplexingness and edit ineffectiveness. This means that when a model finds new knowledge very perplexing, it is difficult to incorporate this knowledge into the model. Similarly, a person might be resistant to learning something they find hard to understand.
Furthermore, the analysis reveals that certain scenarios exhibit high Pearson coefficients, such as the application of MEMIT to the GPT-2 large model. This variance could stem from the possibility that different models encode perplexingness in distinct manners and that editing methods may interact with this perplexingness uniquely.
@@
6.2 Significantly higher perplexingness of higher hierarchy level knowledge
Do hierarchical relations affect perplexingness? To investigate the effect of hierarchical relations on “perplexingness,” we analyze these two groups in HierarchyData: hypernyms (abstract concepts) and hyponyms (specific concepts). The box plots are included in Appendix D. We conduct $t$-tests for two independent samples to determine if the mean perplexingness of the specific level is statistically lower than that of the abstract level. The results of the $t$-tests are detailed in Table 2, with all values demonstrating statistical significance. Our findings indicate that knowledge on a higher hierarchical level (more abstract) is associated with greater perplexingness for the models. This suggests that hierarchical relations are a factor affecting knowledge perplexingness for language models.
@@ 6.2 Significantly higher perplexingness of higher hierarchy level knowledge
$0.00728^{*}$ $0.00605^{*}$ $1.33e-06^{*}$
Table 2: Comparative analysis of perplexingness in HierarchyData: $t$-test results for specific vs. abstract level distributions (* indicates corresponding entry has $p$-value below 0.05).
Correlations between perplexingness and edit ineffectiveness in HierarchyData
Next, we aim to determine if the correlation between perplexingness and edit ineffectiveness also holds for the HierarchyData dataset. We employ the same method to analyze HierarchyData as analyzing CounterFact, focusing on the Pearson correlation coefficient between perplexingness and edit ineffectiveness. The right panel of Figure 3 provides one of the scatter plots (see Appendix B for other plots), showcasing the application of MEMIT on GPT2-XL. We also calculate the Pearson coefficients, with the results presented in Table 3. In this table, $p$-values below 0.05 are marked with , indicating statistical significance. Our analysis reveals a consistent trend: an increase in perplexingness correlates with poorer efficacy of edits (higher negative log conditional probability). This pattern holds true across all scenarios, except when applying the ROME and MEMIT techniques to the GPT-J model.
@@
Table 3: HierarchyData Pearson correlation between perplexingness and edit ineffectiveness ( indicates corresponding entry has $p$-value below 0.05)
6.3 Relationships between hierarchical relations and edit ineffectiveness
Additionally, we want to determine if hierarchical relations within the knowledge ultimately affect the edit ineffectiveness. Box plots (see Appendix C) are constructed to visually compare the ineffectiveness across the two hierarchical levels. Figure 4 shows one of the examples. Furthermore, we conduct $t$-tests on two independent samples to determine whether the mean of the specific level distribution is significantly lower than that of the abstract level distribution. The $p$-values obtained are documented in Table 4. This finding underscores a markedly lower efficacy in editing knowledge at higher hierarchical levels (more abstract knowledge). Significantly, this discrepancy indicates that hierarchical relationships profoundly affect the efficacy of specific editing techniques, like ROME and MEMIT, when applied to particular models, such as GPT2-Large and GPT2-XL. For fine-tuning and LoRA, the results do not appear to be significant, possibly because these methods can address knowledge at different hierarchical levels similarly. But, how about GPT-J?
Figure 5: Same knowledge perplexingness in different models (HierarchyData). Each line represents a piece of knowledge from HierarchyData, sorted by perplexingness in the GPT-2L model. We observe that GPT-J appears darker in the heatmap, suggesting it finds the same knowledge less perplexing.
@@ 6.3 Relationships between hierarchical relations and edit ineffectiveness
From the previous experiment, we observe that GPT-J did not show any difference in edit efficacy when editing higher hierarchy and lower hierarchy knowledge. To determine if GPT-J finds the same knowledge less perplexing compared to GPT-2L and GPT-2XL, we generated a heatmap of each knowledge’s perplexingness in the HierarchyData for each model, as shown in Figure 5. Each line represents a piece of knowledge in the HierarchyData, sorted by perplexingness in the GPT-2L model. We observed that GPT-J appears darker in the heatmap, indicating it finds the same knowledge less perplexing.
@@
Table 4: Comparative analysis of ineffectiveness in HierarchyData: $t$-test results for specific vs. abstract level distributions (* indicates corresponding entry has $p$-value below 0.05).
@@ 7 Discussion
Our experimental results reveal intriguing variations in how different models handle perplexing knowledge, particularly in the context of editing. Specifically, the application of ROME and MEMIT to GPT-J exhibits a notably low Pearson correlation between perplexingness and editing ineffectiveness. Moreover, within the HierarchyData context, these correlations appear insignificant. Additionally, the influence of hierarchical relations on the editing ineffectiveness of ROME and MEMIT when applied to GPT-J seems negligible. This suggests that GPT-J may employ a unique mechanism for storing and processing different hierarchy-level knowledge compared to other models. These differences highlight the need to comprehend each model’s unique architecture and methods for handling perplexing concepts, suggesting a move towards tailored editing strategies.
@@
An intuition is that when editing towards a hypernym (“animal” $\rightarrow$ “plant”), it is assumed that the hyponym (“cat” $\rightarrow$ “plant”) is edited as well, making the edit of hypernym inherently harder. Yet, the dependent knowledge is usually not edited, for popular editing methods (Li et al., 2023b).
Are there other factors that may influence the perplexingness?
The investigation into the responsiveness of different editing techniques to perplexing knowledge reveals that FT and LoRA are seemingly unaffected by the hierarchical structure of knowledge. Notably, there exists a pronounced correlation between perplexingness and the ineffectiveness of edits. This suggests that while FT and LoRA are adept at navigating the hierarchical relationships among words, they falter when addressing the inherent perplexingness present within the knowledge. This observation leads to the hypothesis that additional factors, beyond hierarchical complexity, play a pivotal role in influencing perplexingness when employing FT and LoRA for knowledge editing.
Figure 6: Factors that influence the edit efficacy. In this paper, we explore how hierarchical relations influence perplexingness and how perplexingness, in turn, affects the efficacy of edits. However, other factors may also impact the efficacy of edits, such as the choice of edit methods or models. Additionally, there may be other factors contributing to perplexingness, such as fine-grained or ambiguous knowledge. Investigating these aspects is beyond the scope of this work and will be addressed in future research.
@@
The impact of perplexingness on the ineffectiveness of various editing methodologies can vary significantly. Moreover, the manner in which different models interpret, process, and encode the perplexingness of knowledge also differs. This suggests a complex interplay between the editing methods used and the intrinsic mechanisms of the models, as illustrated by Figure 6, underscoring the need for a nuanced understanding of both to optimize knowledge editing strategies. Other factors beyond hierarchical complexity may also contribute to the perplexingness of knowledge. For instance, knowledge with significant semantic overlap with other concepts can introduce perplexingness by creating competing or conflicting representations. Similarly, fine-grained or highly specific knowledge, as well as ambiguous knowledge with multiple possible interpretations, may further increase perplexingness.
@@
Future model editing efforts should pay attention to understanding the nature of the knowledge being edited, particularly its level of perplexingness. To aid in this endeavor, we have introduced a hierarchy dataset designed to facilitate it. It is crucial to ensure that editing methods are versatile and effective across a diverse range of data types. Moreover, adopting different editing approaches tailored to the specificities of each model can significantly enhance the success of edits. When editing hierarchy knowledge, we can try to use edit methods like fine-tuning or LoRA. It may dismiss the influence of hierarchy data. Also, we should pay attention to the side effects of knowledge edit.
@@ 8 Conclusion
Our investigation into knowledge editing in LLMs reveals a fundamental challenge: the more perplexing a piece of knowledge is to an LLM, the more resistant it becomes to modification through existing editing methods. Through comprehensive analysis using both the CounterFact dataset and our newly developed HierarchyData, we demonstrate that abstract concepts (hypernyms) are inherently more perplexing to LLMs than their specific counterparts (hyponyms), leading to lower editing efficacy. These findings not only highlight a previously unexplored aspect of model editing technology but also provide crucial insights for developing more sophisticated editing methodologies that can effectively handle knowledge across different levels of conceptual abstraction.
@@ 9 Limitation
In this paper, we focus on a short hierarchy chain to facilitate the comparison between higher and lower hierarchy levels. Future works can explore longer hierarchy chains. The experiment can be scaled up, including the use of larger models and larger datasets. Additional types of evaluation can be applied. For instance, we could ask language model-specific questions to determine if the knowledge has actually been edited. However, this approach is very labor-intensive and was not implemented in this study.
Appendix A Correlation of perplexingness and efficacy in CounterFact
We plot the perplexingness (pre-edit probabilities of the new target) against the efficacy (post-edit probabilities of the new target) to visually analyze their relationship. This analysis is conducted using the first 2000 groupings from the CounterFact dataset. Figure 7 displays the scatter plot for editing methods applied to GPT2-Large. Similarly, Figure 8 presents the scatter plot for methods used on GPT2-XL, and Figure 9 illustrates the scatter plot for edits performed on GPT-J(6B).
Figure 7: Pre vs. post probability of new knowledge (CounterFact) on GPT2-Large using a. FT (first) b. LoRA (second) c. ROME (third) d. MEMIT (fourth).
Figure 8: Pre vs. post probability of new knowledge (CounterFact) on GPT2-XL using a. FT (first) b. LoRA (second) c. ROME (third) d. MEMIT (fourth).
Figure 9: Pre vs. post probability of new knowledge (CounterFact) on GPT-J(6B) using a. FT (first) b. LoRA (second) c. ROME (third) d. MEMIT (fourth).
Appendix B Correlation of perplexingness and efficacy in HierarchyData
To visually explore the relationship between perplexingness and editing efficacy, we plot these dimensions against each other using 198 groupings from the HierarchyData dataset. Figure 10 shows the scatter plot highlighting the effects of editing methods on the GPT2-Large model. Likewise, Figure 11 demonstrates the scatter plot for the GPT2-XL model, and Figure 12 displays the scatter plot for edits on the GPT-J(6B) model, providing a clear visual representation of how perplexingness correlates with the efficacy of knowledge edits across different models.
Figure 10: Pre vs. post probability of new knowledge (HierarchyData) on GPT2-Large using a. FT (first) b. LoRA (second) c. ROME (third) d. MEMIT (fourth).
Figure 11: Pre vs. post probability of new knowledge (HierarchyData) on GPT2-XL using a. FT (first) b. LoRA (second) c. ROME (third) d. MEMIT (fourth).
Figure 12: Pre vs. post probability of new knowledge (HierarchyData) on GPT-J(6B) using a. FT (first) b. LoRA (second) c. ROME (third) d. MEMIT (fourth).
@@ Appendix B Correlation of perplexingness and efficacy in HierarchyData
Figure 13: Specific vs. abstract probability distribution (HierarchyData) on GPT2-Large using a. FT (first) b. LoRA (second) c. ROME (third) d. MEMIT (fourth).
Figure 14: Specific vs. abstract probability distribution (HierarchyData) on GPT2-XL using a. FT (first) b. LoRA (second) c. ROME (third) d. MEMIT (fourth).
Figure 15: Specific vs. abstract probability distribution (HierarchyData) on GPT-J(6B) using a. FT (first) b. LoRA (second) c. ROME (third) d. MEMIT (fourth).
@@
We perform a comparative analysis of the perplexingness across both specific and abstract hierarchical levels by plotting their distributions. This analysis is based on 198 instances from the HierarchyData dataset, evenly divided between 99 specific-level and 99 abstract-level cases. Figure 16 presents the box plots, illustrating the impact of editing methods on the GPT2-Large, GPT2-XL, and GPT-J(6B) models, thereby offering insights into the variation of perplexingness across different levels of hierarchy and models.
Figure 16: Pre-edit specific vs. abstract probability distribution (HierarchyData) on a. GPT2-Large (first) b. GPT2-XL (second) c. GPT-J(6B) (third).